The voltage across a 10 μF capacitor to charge it with 10 mC is

1
Expert's answer
2022-01-24T18:45:30-0500

C = 10 uF = "10 * 10^{-6} F"

Q = 10 mC = "10 * 10^{-3} C"


Q = C V

Therefore

V = "\frac{Q}{C}" = "\frac{10*10^{-3}}{10*10^{-6}}" = 1000V


V = 1000 V